Ammonium perchlorate NH

4

ClO

4

 is the solid rocket fuel used by the U.S. Space Shuttle. It reacts with itself to produce nitrogen gas N

2

, chlorine gas Cl

2

, oxygen gas O

2

, water H

2

O

, and a great deal of energy.

What mass of chlorine gas is produced by the reaction of 

7.59

g


of ammonium perchlorate?


1
Expert's answer
2021-09-16T02:01:47-0400

2NH4ClO4 → N2 + Cl2 + 2O2 + 4H2O

M(NH4ClO4) = 117.5 g/mol

"n = \frac{m}{M}"

n(NH4ClO4) "= \frac{7.59}{117.5}=0.06459 \;mol"

According to the reaction equation

n(Cl2) "= \frac{1}{2}" n(NH4ClO4)

"= \frac{1}{2} \times 0.06459 \\

= 0.03229 \;mol"

M(Cl2) = 73.93 g/mol

m(Cl2) "= 0.03229 \times 73.93 = 2.38 \;g"

Answer: 2.38 g